OpenStudy (anonymous):

lim t->0 t^3 /tan^3(2t) ????

OpenStudy (anonymous):

Near 0z we have tan x behave like x so our limit behaves like t^3/(8t^3) = 1/8

OpenStudy (anonymous):

as \(t\to0\) we observe \(\tan 2t\approx2t\) and therefore \(\tan^32t\approx(2t)^3=8t^3\)

OpenStudy (primeralph):